# ORMNet: Object-centric Relationship Modeling for Egocentric Hand-object Segmentation

### v1

@@ III-B Overall Architecture
The Overall diagram of the proposed ORMNet is shown in Fig. 2. The model consists of an encoder, a bottleneck, and a decoder with multiple branches. The encoder is used to extract distinguishing features of the input egocentric image, the bottleneck is used to enhance the extracted global features, and different branches of the decoder are used to predict the final segmentation masks for diverse categories. Following the previous work[4], ORMNet also contains the contact boundary (CB) decoder branch to predict the mask of the contact boundary. For simplification, we remove the CB decoder branch in Fig. 2.
III-C Encoder and Bottleneck
@@
Taking an egocentric image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ as input, the first step of ORMNet is to use the image encoder to extract the multi-scale features $\textbf{F}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ by a series of down-sampling by a factor of 2, where $H_{i}$, $W_{i}$, and $C_{i}$ are the height, width, and channel of the $i^{th}$ stage feature map, respectively. The image encoder used in practice is the Swin Transformer encoder[41]. The extraction process of generating multi-scale features in the encoder can be expressed as follows:
$\mathbf{F}_{i+1}=f_{enc}^{i+1}(\mathbf{F_{i}}|W_{enc}^{i+1}),i\in\left\{0,1,2\right\},$ (1)
@@
where the $f_{enc}^{i}(\cdot)$ denotes the encoder stage with the down-sampling for generating the $i^{th}$ stage feature map, and the $W_{enc}^{i}$ represents corresponding parameters.
@@
After generating the multi-scale features, the final feature map with the smallest height and width and the largest channel is sent into the bottleneck to enhance the representation of the encoded feature. Following the work[15], the bottleneck is composed of a patch merging layer and two Swin Transformer blocks , which is denoted as follows:
$\mathbf{F}_{ehc}=f_{bot}(\mathbf{F}_{last}|W_{bot}),$ (3)
where the $\mathbf{F}_{last}$ and $\mathbf{F}_{ehc}$ represent the final encoded feature map and enhanced final feature map, and the $f_{bot}(\cdot)$ denotes the bottleneck network with the parameter of $W_{bot}$. The obtained enhanced memory feature $\mathbf{F}_{ehc}$ and multi-scale features $\{\mathbf{F}_{i}|i=0,1,2\}$ are then passed into separate decoder branches to predict the hands and interacting objects.
III-D Hand-Object Relation Module
@@
Figure 3: Detailed architecture of the proposed HOR module. Taking the features of hands and objects as input, this module accomplishes hand-guided attention to model the relationship between the hands and interacting objects, promoting the object feature learning of the network.
@@
Among the egocentric visual information, the hand information of the HMD wearer is the most prominent and representative. Moreover, the interacted objects can only be identified based on whether they are in contact with the hand rather than appearance and texture cues. Therefore, using hand features to enhance the extraction and expression of object features is beneficial to the EgoHOS task, which has been ignored in previous work. This paper proposes the HOR module to model the correlation between hand features and object features in interaction. Specifically, this module uses representative hand features as priors and guidance and utilizes the hand-guided attention mechanism to guide the learning of object features, as shown in Fig. 3.
@@
Taking the hand features $\textbf{F}_{hand}\in\mathbb{R}^{\hat{C}\times\hat{H}\times\hat{W}}$ and the object features $\textbf{F}_{obj}\in\mathbb{R}^{\hat{C}\times\hat{H}\times\hat{W}}$ generated by decoder branches as the input, this module aims to use hand features to enhance the object features to focus more precisely on the interacting region, where $\hat{C}$, $\hat{H}$, and $\hat{W}$ are the channel, height, and width of the feature map. First, two stacked convolution operations are used after the LayerNorm layer to generate the query, key, and value. The query $Q\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ is generated from hand features, and the key $K\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ and value $V\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ are generated by object features, which can be denoted as follows,
$Q=f_{conv}(f_{conv}(f_{ln}(\textbf{F}_{hand}))),$ (4)
$K,V=f_{conv}(f_{conv}(f_{ln}(\textbf{F}_{obj}))).\vspace{-0.3cm}$ (5)
@@
After that, the query and key are used to perform the hand-guided attention operation, i.e.,
$\textbf{A}=\delta(Matmul(Norm(Q),Norm(K^{T}))),$ (6)
$\textbf{F}_{o}=Conv(Matmul(\textbf{A},V)),$ (7)
@@
where the $\textbf{F}_{o}$ is the output after convolution, the $Matnul$ means the matrix multiplication, and the $\delta(\cdot)$ represents the softmax function. Then, the input hand features are added to the $\textbf{F}_{o}$ to generate the guided features. Finally, the addition between guided features and original object features is performed to enhance the representation of the object features, which can be described as follows,
$\textbf{F}_{guide}=\textbf{F}_{o}+\textbf{F}_{hand},\quad\hat{\textbf{F}}_{obj}=\textbf{F}_{obj}+\textbf{F}_{guide},$ (8)
where the $\hat{\textbf{F}}_{obj}$ is the output enhanced object features.
III-E Object Relation Decoupling Strategy
@@
Based on the proposed HOR module, a naive way is to use one decoder branch to predict hands and another decoder branch to predict interacting objects. Then, the HOR module can establish the relation between hands and objects, as shown in Sec. III-D.
However, this paper argues that in this way, the three different types of objects, i.e., left-hand objects, right-hand objects, and two-hand objects, are coupled, which will cause confusion and ambiguity when training the network. In particular, if the network is trained to predict left-hand objects and right-hand objects based on whether they are in contact with the left and right hands, then two-hand objects that are in contact with both hands simultaneously are likely to be confused with either left-hand objects or right-hand objects.
Therefore, different from previous work, our method proposes the ORD strategy to decouple the three kinds of objects according to their physical meanings. Specifically, we consider two-hand objects to be both left-hand and right-hand objects so that two-hand objects can be represented by the intersection of the other two categories. This approach establishes a relationship between two-hand objects and the other two categories that are originally unrelated, enabling the network to concentrate solely on the hand-object interaction without classifying the category.
In practice, the ORD process can be divided into data preparation, training, and inference. First, in the phase of data preparation, we consider the ground truth (GT) of two-hand objects in the dataset as both left-hand objects and right-hand objects, so we generate the new GT of left- and right-hand objects by union operation, as shown in Fig. 2 (left). For example, if the value in the two-hand object GT $\textbf{G}_{to}$ at the position $(i,j)$ is 1, then the value of the left-hand object GT $\textbf{G}_{lo}$ and right-hand object GT $\textbf{G}_{ro}$ at $(i,j)$ will be set to 1. The process is shown below,
$\hat{\textbf{G}}_{lo}=\textbf{G}_{lo}\bigcup\textbf{G}_{to},\quad\hat{\textbf{G}}_{ro}=\textbf{G}_{ro}\bigcup\textbf{G}_{to},$ (9)
@@
where the $\bigcup$ means the union operation, the $\hat{\textbf{G}}_{lo}$ and $\hat{\textbf{G}}_{ro}$ are the generated GT labels for left-hand objects and right-hand objects.
Then, in the training process, we only use $\hat{\textbf{G}}_{lo}$ and $\hat{\textbf{G}}_{ro}$ to supervise the two object prediction branches, as shown in Fig. 2, which can be expressed as follows,
$\hat{\textbf{M}}_{O}^{L}=f_{dec}^{lo}(\textbf{F}_{ehc},\textbf{F},W_{dec}^{lo}),\textbf{F}=\{\textbf{F}_{i}|i=0,1,2\},$ (10)
$\hat{\textbf{M}}_{O}^{R}=f_{dec}^{ro}(\textbf{F}_{ehc},\textbf{F},W_{dec}^{ro}),\textbf{F}=\{\textbf{F}_{i}|i=0,1,2\},$ (11)
$\mathscr{L}_{lo}=l_{ce}(\hat{\textbf{M}}_{O}^{L},\hat{\textbf{G}}_{lo}),\quad\mathscr{L}_{ro}=l_{ce}(\hat{\textbf{M}}_{O}^{R},\hat{\textbf{G}}_{ro})),$ (12)
@@
where the $\hat{\textbf{M}}_{O}^{L}$ and $\hat{\textbf{M}}_{O}^{R}$ are the predicted left-hand objects and right-hand objects. The $f_{dec}^{lo}(\cdot)$ and $f_{dec}^{ro}(\cdot)$ represent the decoder branch of left-hand objects and right-hand objects with parameters of $W_{dec}^{lo}$ and $W_{dec}^{ro}$. And the $\mathscr{L}_{lo}$ and $\mathscr{L}_{ro}$ denote the loss function of left-hand objects and right-hand objects, which we use the cross entropy $l_{ce}(\cdot)$ in practice.
@@
In inference, based on the predicted left- and right-hand objects $\hat{\textbf{M}}_{O}^{L}$ and $\hat{\textbf{M}}_{O}^{R}$, the two-hand objects are generated by a simple but effective operation, i.e., intersection, as shown in Fig. 2 (right) and follows,
$\textbf{M}_{O}^{T}=\hat{\textbf{M}}_{O}^{L}\bigcap\hat{\textbf{M}}_{O}^{R},$ (13)
$\textbf{M}_{O}^{L}=\hat{\textbf{M}}_{O}^{L}-\textbf{M}_{O}^{T},\quad\textbf{M}_{O}^{R}=\hat{\textbf{M}}_{O}^{R}-\textbf{M}_{O}^{T},$
@@
where the $\bigcap$ means the intersection operation, and the $\textbf{M}_{O}^{L}$ and $\textbf{M}_{O}^{R}$ are the final left- and right-hand object predictions.
The proposed ORD strategy implicates the physical meaning of the three kinds of objects, especially the meaning of the two-hand objects, i.e., objects that interact with both hands simultaneously. Furthermore, in this way, the network can be trained only to predict the objects that interact with the left-/right-hand rather than focusing on classifying the category of the objects, which achieves decoupling the objects and decreasing the confusion and ambiguity of the algorithm.
III-F Training and inference
@@
The overall ORMNet is trained by the weighted cross-entropy loss function between predictions and GTs of hands, left-hand objects, right-hand objects, and contact boundaries. The overall loss function $\mathscr{L}$ can be computed by,
$\mathscr{L}=\alpha(l_{ce}(\textbf{M}_{O}^{L},\hat{\textbf{G}}_{lo})+l_{ce}(\textbf{M}_{O}^{R},\hat{\textbf{G}}_{ro}),)+$ (14)
@@
where the $l_{ce}(x,y)$ means the cross entropy loss between $x$ and $y$, the $\textbf{M}_{CB}$ and $\textbf{G}_{cb}$ are the prediction and GT of the contact boundary, respectively. The GT of the contact boundary is generated by the GT of the hands and interacting objects automatically, which is the overlapping region of the dilated hand and object masks [4].
For inference, as stated in Sec. III-E, the network predicts the masks of hands, left-hand objects, and right-hand objects, and then the mask of the two-hand objects can be obtained by the proposed ORD strategy.
TABLE I: Comparison results on the EgoHOS in-domain test set measured by IoU/Acc and mIoU/mAcc. The ``"`$\prime$"` ` ′ " means the method with data augmentation, the ``"``"`$\diamondsuit$"` ` ♢ " means the method introduced contact boundary, and the ``"``"`$\flat$"` ` ♭ " means the method with data augmentation and contact boundary.

### v2

@@ 3.2 Overall Architecture
The Overall diagram of the proposed ORMNet is shown in Fig. 2, which consists of data preparation (left), training (middle), and inference (right). The detailed information about data preparation and inference will be introduced in Section 3.4. The model architecture consists of an encoder, a bottleneck, and a decoder with multiple branches. The encoder is used to extract multi-scale distinguishing features of the input egocentric image, the bottleneck is used to enhance the extracted global features, and different branches of the decoder are used to predict the final segmentation masks for diverse categories.
@@
Specifically, taking an egocentric image $\mathbf{I}$ as input, the first step of ORMNet is to use the image encoder to extract the multi-scale representative features $\textbf{F}=\{\textbf{F}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}\;|\;i=0,1,\cdots,N\}$ by a series of down-samplings with a factor of 2, where $H_{i}$, $W_{i}$, and $C_{i}$ are the height, width, and channel of the $i^{th}$-stage feature map, respectively. The image encoder used in practice is the Swin Transformer encoder [41], and N is set to 3. The extraction process of generating multi-scale features in the encoder can be expressed as follows:
$\mathbf{F}_{i+1}=f_{enc}^{i+1}(\mathbf{F_{i}}\;|\;W_{enc}^{i+1}),\;i\in\left\{0,1,\cdots,N-1\right\},$ (1)
@@
where the $f_{enc}^{i}(\cdot)$ denotes the encoder stage for generating the $i^{th}$-stage feature map, and the $W_{enc}^{i}$ represents corresponding parameters.
@@
After generating the multi-scale features, the final global feature map $\textbf{F}_{N}$ is sent into the bottleneck to enhance the representation of the encoded feature. Following the work [18], the bottleneck is composed of a patch merging layer and two Swin Transformer blocks, which are denoted as follows:
$\mathbf{F}_{ehc}=f_{bot}(\mathbf{F}_{N}\;|\;W_{bot}),$ (3)
where the $\mathbf{F}_{ehc}$ represents the enhanced final feature map, and the $f_{bot}(\cdot)$ denotes the bottleneck network with the parameter of $W_{bot}$.
Taking the enhanced global feature $\mathbf{F}_{ehc}$ and multi-scale features $\textbf{F}^{\prime}=\{\textbf{F}_{i}\;|\;i=0,1,\cdots,N-1\}$ as input, different decoder branches are used to predict the final mask of each category. Following the previous work [9], ORMNet also contains the contact boundary (CB) decoder branch to predict the mask of the contact boundary. For simplification, we remove it in Fig. 2 (middle). In each decoder branch, N+1 decoder stages are employed to predict the masks. And the output feature of each stage is concatenated with corresponding encoded features by skip-connection [42] as the input to the next stage. The final up-sampling layer is used to recover the resolution to the same as the input egocentric image. The process of each decoder branch can be denoted as:
$\textbf{D}_{0}=f_{up}^{0}(\textbf{F}_{ehc}),$ (4)
$\textbf{D}_{i+1}=f_{up}^{i+1}(f_{cat}(\textbf{D}_{i},\textbf{F}_{N-1-i})),$ (5)
$\textbf{M}=f_{us}(\textbf{D}_{N}),$ (6)
where the $f_{up}^{i}(\cdot)$ means the $i^{th}$ decoder stage, which outputs $\textbf{D}_{i}$ as decoded feature map. The $f_{cat}(\cdot)$ means the concatenation operation, the $f_{us}(\cdot)$ represents the final up-sampling layer, and the M denotes the predicted mask of this decoder branch.
@@
Figure 3: Detailed architecture of the proposed HOR module. Taking the features of hands and interacting objects as input, this module accomplishes hand-guided attention to model the relationship between them, promoting the object feature learning of the network.
3.3 Hand-Object Relation Module
@@
Among the egocentric visual information, the hand information of the HMD wearer is the most prominent and representative because we can easily recognize hands by shape and texture [20]. Moreover, the contacted objects can be accurately identified based on whether they are interacting with hands rather than appearance and texture cues. Therefore, using hand features to enhance the extraction and expression of interacting-object features is beneficial to the EgoHOS task. While certain sequential methods [9] employ predicted hand masks to segment interacting objects, the performance of this coarse approach cannot effectively excavate the relationships between them. This paper proposes the HOR module to model the correlation between hand and object features in interaction, which uses representative hand features as priors and guidance and utilizes the hand-guided attention mechanism to guide the learning of interacting-object features, as shown in Fig. 3.
@@
Assume we use separate decoder branches for hands and interacting objects, taking the $N^{th}$ layer of hand features $\textbf{D}_{hand}\in\mathbb{R}^{\hat{C}\times\hat{H}\times\hat{W}}$ from the hand decoder branch and the object features $\textbf{D}_{obj}\in\mathbb{R}^{\hat{C}\times\hat{H}\times\hat{W}}$ generated by the object decoder branch as input, this module aims to use hand features to enhance the representation of interacting objects, where $\hat{C}$, $\hat{H}$, and $\hat{W}$ are the channel, height, and width of the feature map. First, two stacked convolution operations are used after the LayerNorm layer to generate the query, key, and value. The query $Q\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ is generated from hand features, and the key $K\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ and value $V\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ are generated by interacting-object features, which can be denoted as follows:
$Q=f_{conv}(f_{conv}(f_{ln}(\textbf{D}_{hand}))),$ (7)
$K,V=f_{conv}(f_{conv}(f_{ln}(\textbf{D}_{obj}))),$ (8)
@@
where the $f_{conv}(\cdot)$ and $f_{ln}(\cdot)$ represent the convolution and LayerNorm layers. After that, the query, key, and value are used to perform the hand-guided attention operation, i.e.:
$\textbf{A}=\delta(f_{Matmul}(Norm(Q),Norm(K^{T}))),$ (9)
$\textbf{F}_{o}=f_{conv}(f_{Matmul}(\textbf{A},V)),$ (10)
@@
where the $\textbf{F}_{o}$ is the output after convolution, the $f_{Matnul}(x,y)$ means the matrix multiplication between x and y, and the $\delta(\cdot)$ represents the softmax function. Then, the input hand features are added to the $\textbf{F}_{o}$ to generate the guided features. Finally, the element-wise addition between guided features and original object features is performed to enhance the representation of the interacting-object features, which can be described as follows:
$\textbf{F}_{guide}=\textbf{F}_{o}+\textbf{D}_{hand},$ (11)
${\textbf{D}}_{obj}^{\prime}=\textbf{D}_{obj}+\textbf{F}_{guide},$ (12)
where the ${\textbf{D}}_{obj}^{\prime}$ is the final enhanced interacting-object features. In this way, the relationships between hands and interacting objects are established explicitly by the hand-guided attention mechanism. Therefore, hand features can be used as queries to extract more representative and distinguishing object features, improving the interacting-object segmentation performance.
3.4 Object Relation Decoupling Strategy
@@
Based on the proposed HOR module, a naive way is to use one decoder branch to predict hands and another decoder branch to predict interacting objects. Then, the HOR module can establish the relation between hands and interacting objects, as shown in Sec. 3.3. However, this paper argues that in this way, the three different types of interacting objects, i.e., left-hand objects, right-hand objects, and two-hand objects, are coupled.
To be specific, we observes that the two-hand object interacts with both left and right hands, so it encompasses attributes of both the left-hand and right-hand objects. Based on this observation, we contend that two-hand objects can be conceptualized as the intersection of the remaining two categories (left-hand object and right-hand object). Therefore, if we remove the two-hand object category during training and represent this category by using the other two categories in inference, the network can be relieved from classifying “whether the object is interacting with both hands at the same time” and focus only on the left- and right-interaction.
Therefore, different from previous work, our method proposes the ORD strategy to decouple the three kinds of interacting objects according to observed relations. In practice, the ORD process can be divided into data preparation, training, and inference. During data preparation, since two-hand objects encompass attributes of left-hand and right-hand objects, the ground truth (GT) of the two-hand object is treated as the left-hand and right-hand objects to generate new object labels. Specifically, as shown in Fig. 2 (left), the new GT of the left/right-hand object is obtained by the union operation between the original left/right-hand object and two-hand object labels. The process is shown below:
${\textbf{G}}_{lo}^{\prime}=\textbf{G}_{lo}\cup\textbf{G}_{to}=\{(x\in\textbf{G}_{lo})\vee(x\in\textbf{G}_{to})\},$ (13)
${\textbf{G}}_{ro}^{\prime}=\textbf{G}_{ro}\cup\textbf{G}_{to}=\{(x\in\textbf{G}_{ro})\vee(x\in\textbf{G}_{to})\},$ (14)
@@
where the $\cup$ means the union operation, the ${\textbf{G}}_{lo}$, ${\textbf{G}}_{ro}$, and ${\textbf{G}}_{to}$ are the original labels for left-hand objects, right-hand objects, and two-hand objects, respectively. The ${\textbf{G}}_{lo}^{\prime}$ and ${\textbf{G}}_{ro}^{\prime}$ represent the generated new supervision signals for left-hand objects and right-hand objects.
Then, during training, we use separate decoder branches to predict the left-hand objects and right-hand objects, as shown in Fig. 2 (middle). The left-hand object decoder output $\textbf{M}_{lo}$ can be obtained by up-sampling the outcome $\textbf{D}_{lo}^{\prime}$ of the HOR module from equation 12, and the right-hand object decoder output $\textbf{M}_{ro}$ can be obtained by up-sampling the outcome $\textbf{D}_{ro}^{\prime}$ of the HOR module from equation 12. ${\textbf{G}}_{lo}^{\prime}$ and ${\textbf{G}}_{ro}^{\prime}$ are utilized to supervise these two branches, which can be expressed as follows,
$\mathscr{L}_{lo}=l_{ce}(\textbf{M}_{lo},\textbf{G}_{lo}^{\prime}),$ (15)
$\mathscr{L}_{ro}=l_{ce}(\textbf{M}_{ro},\textbf{G}_{ro}^{\prime}),$ (16)
@@
where the $\mathscr{L}_{lo}$ and $\mathscr{L}_{ro}$ denote the loss of left-hand objects and right-hand objects, which we use the cross entropy $l_{ce}(\cdot)$ in practice.
@@
In inference, after obtaining the decoder output of left- and right-hand objects ${\textbf{M}}_{lo}$ and ${\textbf{M}}_{ro}$, the results of two-hand objects are generated by an ingenious operation, i.e., intersection. The final predictions of the left-hand objects and the right-hand objects have to remove the region of the two-hand objects, as shown in Fig. 2 (right) and follows,
$\textbf{M}_{O}^{T}={\textbf{M}}_{lo}\cap{\textbf{M}}_{ro}=\{(x\in\textbf{M}_{lo})\wedge(x\in\textbf{M}_{ro})\},$ (17)
${\textbf{M}}_{O}^{L}={\textbf{M}}_{lo}-\textbf{M}_{O}^{T}=\{(x\in\textbf{M}_{lo})\wedge(x\notin\textbf{M}_{O}^{T})\},$ (18)
${\textbf{M}}_{O}^{R}={\textbf{M}}_{ro}-\textbf{M}_{O}^{T}=\{(x\in\textbf{M}_{ro})\wedge(x\notin\textbf{M}_{O}^{T})\},$ (19)
@@
where the $\cap$ means the intersection operation. The ${\textbf{M}}_{O}^{L}$, ${\textbf{M}}_{O}^{R}$, and ${\textbf{M}}_{O}^{T}$ are the final left-, right-, and two-hand object predictions, respectively.
The innovative ORD strategy decouples the two-hand objects during training using pure mathematical intersection and union operations. This approach enables the algorithm to prioritize the interaction between hands and objects, thereby diminishing the emphasis on classification and decreasing the confusion and ambiguity of the algorithm.
3.5 Training and Inference
@@
The overall ORMNet is trained by the weighted cross-entropy loss function between predictions and GTs of hands, left-hand objects, right-hand objects, and contact boundaries [9]. The overall loss function $\mathscr{L}$ can be computed by:
$\mathscr{L}=\alpha(l_{ce}(\textbf{M}_{lo},{\textbf{G}}_{lo}^{\prime})+l_{ce}(\textbf{M}_{ro},{\textbf{G}}_{ro}^{\prime}))+$ (20)
@@
where $\alpha,\;\gamma,$ and $\lambda$ denote the loss weight of interacting objects, hands, and contact boundaries [9], respectively. $\textbf{M}_{CB}$ presents the output of the contact boundary decoder branch, which is supervised by $\textbf{G}_{cb}$. Following previous work [9], the GT of the contact boundary is generated by the overlapping region of the dilated hand and interacting-object masks. $\textbf{M}_{H}$ denotes the output of the hand decoder branch, which is supervised by $\textbf{G}_{h}$.
For inference, as stated in Sec. 3.4, the decoder branches predict the masks of hands, left-hand objects, and right-hand objects. Then, the final predicted masks of left-hand objects, right-hand objects, and two-hand objects can be obtained by the proposed ORD strategy. The overall diagram for training and inference is shown in Algorithm 1.
Algorithm 1 ORMNet
  ORD strategy - Data Preparation
        ${\textbf{G}}_{lo}^{\prime}=\textbf{G}_{lo}\bigcup\textbf{G}_{to}$
        ${\textbf{G}}_{ro}^{\prime}=\textbf{G}_{ro}\bigcup\textbf{G}_{to}$
  TRAIN
[…]
       $\textbf{Prediction: }{\textbf{M}}_{lo},{\textbf{M}}_{ro},\textbf{M}_{H},\textbf{M}_{CB}$
[…]
       Return: $W$
  Inference
       $\textbf{Prediction: }{\textbf{M}}_{lo},{\textbf{M}}_{ro},\textbf{M}_{H},\textbf{M}_{CB}$
       $\textbf{ORD Strategy - Inference: }\textbf{M}_{O}^{T}={\textbf{M}}_{lo}\bigcap{\textbf{M}}_{ro}$
                          $\textbf{M}_{O}^{L}={\textbf{M}}_{lo}-\textbf{M}_{O}^{T}$
                          $\textbf{M}_{O}^{R}={\textbf{M}}_{ro}-\textbf{M}_{O}^{T}$
       $\textbf{Return: }\textbf{M}_{H},\textbf{M}_{CB},\textbf{M}_{O}^{T},\textbf{M}_{O}^{L},\textbf{M}_{O}^{R}$
4 Experiments
4.1 Datasets and Metrics
To validate the effectiveness of the proposed ORMNet, we compare its performance with different methods on the EgoHOS [9] in-domain test set. Additionally, we employed the EgoHOS out-of-domain test set as well as a dataset we developed from the HOI4D [43] dataset, referred to as mini-HOI4D, to assess the generalization capability of our ORMNet.
EgoHOS. The EgoHOS [9] dataset consists of 11,743 egocentric images containing per-pixel segmentation labels of hands and interacting objects gathered from Ego4D [24], EPICKITCHEN [26], THU-READ [44], and their own collected egocentric videos. Among them, 8,993 images are used for training, 1,124 images are for validation, 1,126 images are used for in-domain testing, and 500 images are used for out-of-domain testing.
mini-HOI4D. Since the EgoHOS task is an emerging problem, to the best of our knowledge, no suitable datasets other than the EgoHOS dataset can be used to validate the generalization of the model. As a result, we selected the relatively similar HOI4D [43] dataset and further processed and aligned its annotations to ensure its applicability for the EgoHOS task requirements. Ultimately, we generated a new dataset, named mini-HOI4D, comprising 1,095 egocentric images with corresponding hands and interacting object annotations, which can be used to assess and compare the generalization capabilities of our ORMNet and other approaches.
This paper uses the most common and widely used metrics in the segmentation task to evaluate the performance, i.e., Intersection-over-Union (IoU) and pixel Accuracy (Acc) for each category. Mean IoU (mIoU) and mean accuracy (mAcc) are also employed to indicate the overall performance.
4.2 Implementation Details
The experiments were conducted on 4 NVIDIA RTX 6000 Ada GPUs. The batch size was set to 12. Data preparation included cropping the image to (448, 448) and normalization with the mean of [106.011, 95.400, 87.429], and the standard variation was set to [64.357, 60.889, 61.419]. Following previous work [9], we also used the Swin Transformer as the backbone. The patch size was set to 4, and the window size was set to 12. The AdamW optimizer was used in the experiment, the learning rate was set to 1e-5, and the weight decay was set to 0.01. The learning rate was set dynamically, i.e., from 0 to 10,000 iterations, the learning rate increased linearly to 1e-4, and from 10,000 iterations to the maximum iteration, the learning rate decreased linearly to 0. The maximum number of iterations in the experiment was set to 180,000.
4.3 Quantitative Results
4.3.1 In-domain Results
Table 1: Comparison results on the EgoHOS in-domain test set measured by IoU/Acc and mIoU/mAcc. The blue letters represent the best results. The ``"`$\prime$"` ` ′ " means the method with data augmentation, the ``"``"`$\diamondsuit$"` ` ♢ " means the method introduced contact boundary, and the ``"``"`$\flat$"` ` ♭ " means the method with data augmentation and contact boundary.